# Fine-tuning Large Language Models for Automated Diagnostic Screening Summaries

### v1

@@ 3.1 MSE questionnaire design
We identified the absence of a standardized MSE questionnaire and reviewed existing options online. We aimed to create a preliminary version tailored to students, encompassing key components like socialness, mood, attention, memory, frustration tolerance, and social support. This process yielded an 18-question questionnaire. Subsequently, we sought the expertise of clinical psychiatrists to refine the questionnaire further. Their valuable insights were instrumental in vetting the relevance and wording of the questions, resulting in a finalized version of the MSE comprising 12 questions. The final MSE questionnaire is provided in the Appendix A.1.
@@ 3.3 Dialogue representation
We developed a Python script to transform participants’ MSE questionnaire responses into simulated doctor-patient conversations to replicate real-world conversations. This process generated 300 doctor-patient conversation sessions, with 3600 (=12 responses x 300 participants) utterances from participants and an equal number from doctors, totaling 7200 utterances. An anonymized excerpt of such conversation for one participant is presented in Table A.2 in the appendix. Figure 2 shows the average length of utterances for each of the 12 questions. The average length of the dialogue conversation with and without the questionnaire is 3662 and 2054 characters, respectively.
@@ 3.4 Reference human summaries
Furthermore, given the structured nature of the MSE questions, the template was well-suited for summarization purposes. The summary template underwent thorough scrutiny through a rigorous review process involving feedback from three independent reviewers (i.e., graduate researchers). Subsequent revisions were made based on their input, ensuring the summary effectively captured key information while maintaining conciseness, clarity, and correctness. After multiple iterations, the final version of the summary template was approved for use by a psychiatrist, leveraging their domain-specific knowledge. The template utilized for creating summaries for each participant can be found in Appendix A.1.
@@ 4 Experiments
The dataset comprising 300 conversations was divided into 200 for training, 50 for validation, and 50 for testing. The Appendix A.2 lists the hyperparameter settings utilized during model training.
@@ 4.1 ROUGE evaluation
We conducted experiments varying the number of epochs for each model to compare their relative performance, as depicted in the Figure A.1 in the Appendix. This figure showcases the models’ adaptability across different ROUGE metrics as epochs increase. Notably, within just five epochs, the ROUGE-1 score of the BART-large-xsum-samsum model surged from 0.290 to 0.736. Similarly, both the BART-base and BART-large-CNN models demonstrated significant improvements in all ROUGE values within the same timeframe. However, the T5 large model failed to exhibit notable adaptation within five epochs.
@@
To gain insight into the model-generated summaries, we conducted experiments with all models across different numbers of epochs (epochs = 5, 10, 25, 50, 100). After analyzing the output summaries generated by these models, we randomly selected one of the participant’s summaries for further analysis. We found that the pre-trained weights of these models tended to produce incomplete summaries, although they were able to capture smaller contexts of the conversation, as shown in Table A.3 in the Appendix.
@@ 4.2.1 Qualitative findings
Moreover, there was a higher degree of contradiction in the generated summary, although a certain level of contradiction was also observed in the human-generated summary. This discrepancy may arise from either the scenario where the domain expert who created the summary may have missed some interpretation or from expert reviewers having different perceptions. Surprisingly, the model did not exhibit hallucination, which is a major problem in large language models. Furthermore, Table A.4 in the Appendix displays the evaluation scores separately from clinicians and non-clinicians. We observed a slight disparity between clinicians and non-clinicians, indicating that clinicians may require a summary with detailed psychological information.
@@ 4.2.2 Inter-rater agreement
We calculated Cohen’s Kappa and Pearson’s correlation coefficient separately for two clinical and three non-clinical annotators (or reviewers). Our clinical annotators achieved a Cohen’s kappa coefficient of 0.45 and a correlation coefficient of 0.87, indicating moderate agreement and strong correlation, respectively. Among non-clinical annotators, annotators 1 and 2 achieved a higher Cohen’s kappa coefficient of 0.7 and a correlation coefficient of 0.97, demonstrating good reliability in their assessments. Table 5(a) displays the Cohen’s Kappa coefficient, while Table 5(b) shows the correlation coefficient among clinical annotators. Similarly, Table 6(a) presents the Cohen’s Kappa coefficient, and Table 6(b) displays the correlation coefficient among non-clinical annotators.
@@ 5 Generalization
Upon reviewing the reviewers’ ratings, we found that the best fine-tuned model’s summary scored well in relevance, fluency, clarity, and repetition as shown in Table 4. However, the generated summary was slightly lacking in terms of missing information, hallucination, and contradiction. Tables A.7 and A.8 in the appendix present dialogue conversations taken from Yao et al. (2022) alongside the corresponding summaries generated by the fine-tuned BART-large-CNN model.
@@ 9 Limitations
Appendix A Appendix
A.1 Summary Template
{spverbatim}
Patient is a ___year old [girl/boy/lady/man]. [His/Her] mood is generally____________[and remains study/but goes up and down] throughout the day. [He/She] [takes/does not take] part in extracurricular activities and _________[socializes/does not socialize] socialize with others. For daily frustration [He/She] (*activities*). [He/She] [feels/does not feel] academic pressure and for this [He/She] (*activities*). [His/Her]concentration and task atten- ding ability is [good/bad]. [He/She] [feels/does not feel] difficulty with memory. [He/She] feels better by (*activities*). [He/She] [feels /does not feel] supported by his family and friends. On a bad day, [he/she] prefers ______________. [He/She] is [experiencing/ not experiencing] _______[stress/anxiety/depression] symptoms such as _____________.
A.2 Hyperparameters
These are the hyperparameters we used across four models - BART base, BART-large-CNN, T5 large, and BART-large-xsum-samsum, using the Pytorch module: { max token length: 1024 tokens, warmup steps: 500, weight decay: 0.01, evaluation strategy: ‘steps’, evaluation steps: 500, save steps: 1e6, gradient accumulation steps: 16 }. The models were trained on an NVIDIA A100-PCIE-40GB GPU, with an average training time of 45 minutes.
MSE Questionnaires
Q1. Please describe your social life at the *anonymized* campus. Are you actively participating in extracurricular activities, interacting with others, or taking initiative to socialize with others?
Q2. Describe your typical daily Mood?
Q3. Does your Mood remain steady or goes up and down throughout the day without any reason or on trivial matters?
Q4. How do you handle day-to-day irritations or frustrations?
Q5. How do you handle pressure related to academics?
Q6. Describe your ability to attend to the task at hand or concentrate on daily tasks (academic, non-academic)?
Q7. Have you noticed any difficulties with memory, such as unable to register new information, forgetting recent events, or not able to recall older personal/factual events?
Q8. What do you do to feel better? For example, some people take caffeine, talk with people, or watch movies to feel better.
Q9. Describe how supported you feel by others (e.g., friends, family) around you and how they help you?
Q10. What do you usually do when you have a bad day or when you are not able to concentrate on work?
Q11. Are you experiencing symptoms of stress, anxiety, or depression? If yes, describe the symptoms?
Q12. Are you doing anything (by self or help seeking) for the ongoing stress, anxiety, or depression, if any? If yes, what?
Table A.1: Finalized MSE Questionnaire
Doctor: What is your patient ID?
Patient: #3A4C307
Doctor: What is your age?
Patient: 19
Doctor: What is your gender?
Patient: Female
Doctor: Please describe your social life at the *anonymized* campus. Are you actively participating in extracurricular activities, interacting with others, or taking initiative to socialize with others?
Patient: I am part of football team and a core member in Physics club and Singularity working team. I also make contacts with my seniors and other staffs for both personal and official discussions. I volunteer for community fests and other initiatives.
Doctor: Describe your typical daily Mood?
Patient: I always try to find happiness in every single moment of my life. But at times I turnout t be moody.
Doctor: Does your Mood remain steady or goes up and down throughout the day without any reason or on trivial matters?
Patient: My mood is dynamic. It goes up and down for both valid and unknown reasons. I get upset on simple jokes and responses from my close circle.
Doctor: How do you handle day-to-day irritations or frustrations?
Patient: I try to connect more with the Almighty through daily prayers. But mostly I prefer sleeping with no disturbance for hours. Nowadays I try to engage myself with a busy schedule and locations.
Doctor: How do you handle pressure related to academics?
Patient: lately I started purposeful ignorance of academic pressure. I will engage my times studying or with close friend. I also try to phone my parents when I feel so exhausted.
Doctor: Describe your ability to attend to the task at hand or concentrate on daily tasks (academic, non-academic)?
Patient: I am mostly able to focus on my task and complete on time. But when I am in a bad mood I will distract myself from the task with social media and resume when I feel fine.
Doctor: Have you noticed any difficulties with memory, such as unable to register new information, forgetting recent events, or not able to recall older personal/factual events?
Patient: Yes I do, and only very lately. I find it very difficult to comprehend what I see and try reading. I also noticed forgetting recent events which where not very important but still to be considered. I also have difficulty in recalling but the least.
Doctor: What do you do to feel better? For example, some people take caffeine, talk with people, or watch movies to feel better.
Patient: Sleep mostly. But if it is with communication gap, I only settle after conveying my last note. I also sing a song or try dancing in my room but I prefer privacy for this
Doctor: Describe how supported you feel by others (e.g., friends, family) around you and how they help you?
Patient: I feel supported very less even from family. And so I don’t expect any support from anyone and try to figure out all alone.
Doctor: What do you usually do when you have a bad day or when you are not able to concentrate on work?
Patient: I sleep for hours or the entire day. I also get some ease after crying or talking about it. I used talk to myself which helped me figure out the situation and motivated to push through.
Doctor: Are you experiencing symptoms of stress, anxiety, or depression? If yes, describe the symptoms?
Patient: Yes, all stress, anxiety and depression
Doctor: Are you doing anything (by self or help seeking) for the ongoing stress, anxiety, or depression, if any? If yes, what?
Patient: Yes, I’m reading books on self-development and self-improvement.
Table A.2: Doctor-patient conversation dialogue of an anonymized participant.
(a) ROUGE-1 score
(b) ROUGE-2 score
(c) ROUGE-L score
(d) ROUGE-L-SUM score
Figure A.1: ROUGE metric obtained after fine-tuning on BART-base, BART-large-CNN, T5 large, and BART-large-xsum-samsum on # epochs= [5,10,25,50,100]
No. of Epochs Generated Summary
0 (Without tuning) Patient: I always try to find happiness in every single moment of my life. I get upset on simple jokes and responses from my close circle. I try to connect more with the Almighty through daily prayers. I sleep for hours or the entire day. I also get some ease after crying or talking about it.
5 Patient is a 19 year old girl. Her mood is generally happy but goes up and down throughout the day. She takes part in extracurricular activities and socializes with others. For daily frustration she tries to connect more with the Almighty. She feels academic pressure. Her concentration and task attending ability is good. She does not feel difficulty with memory. On a bad day, she sleeps for hours, talks to herself, and
10 Patient is a 19 year old girl. Her mood is generally happy but goes up and down throughout the day. She takes part in extracurricular activities and socializes with others. For daily frustration she connects with the Almighty and sleeps. She feels academic pressure. Her concentration and task attending ability is good. She does feel difficulty with memory such as forgetting recent events and older personal/factual events.She feels better by sleeping, singing songs, dancing or dancing in her room. She does not feel supported by her family and friends. On a bad day, she sleeps or talks to herself. She is feeling stress, anxiety and depression symptoms. To overcome this she reads self improvement books.
25 Patient is a 19 year old girl. Her mood is generally happy but goes up and down throughout the day. She takes part in extracurricular activities and socializes with others. For daily frustration she connects with the Almighty and sleeps. She feels academic pressure. Her concentration and task attending ability is okay, she feels difficulty with memory. She does not feel supported by her family and friends. On a bad day, she sleeps or cries and talks to herself. She is feeling stress, anxiety and depression symptoms. To overcome this she reads self development and self improvement books.
50 Patient is a 19 year old girl. Her mood is generally happy but goes up and down throughout the day. She takes part in extracurricular activities and socializes with others. For daily frustration she tries to connect with the lord through daily prayers. She does not feel academic pressure. Her concentration and task attending ability is good. She feels difficulty with memory. She feel better by sleeping, singing songs, dancing, and hiding in her room. She doesn’t feel supported by her family and friends. On a bad day, she sleeps or talks to herself. She is feeling all stress, anxiety and depression symptoms.
100 Patient is a 19 year old girl. Her mood is generally happy but goes up and down throughout the day. She takes part in extracurricular activities and socializes with others. For daily frustration she connects with the Almighty and sleeps. She does not feel academic pressure. Her concentration and task attending ability is good. She feels difficulty with memory such as forgetting recent events and personal facts and recalling the least.. She feels better by sleeping, singing songs, dancing in room, and asking for privacy. She did not feel supported by her family and friends. On a bad day, she sleeps or talks to herself. She is feeling all stress, anxiety and depression symptoms
Table A.3: BART-large-CNN generated summaries at different epochs tested on conversation given in Table A.2 in the Appendix
Reference summary (4.15, 0.48) (4.25, 0.55) (4.15, 0.48) (3.90, 0.64) (1.80, 0.76) (1.15, 0.37) (1.55, 0.60) (1.10, 0.30)
Best model generated summary (3.85, 0.74) (4.10, 0.55) (4.10, 0.55) (4.00, 0.56) (2.20, 0.89) (1, 0) (1.55, 0.75) (1.0, 0.00)
(a) Human evaluation scores obtained by averaging the ratings provided by two clinician on ten conversations
Reference summary (5.00, 0.00) (5.00, 0.00) (4.50, 0.86) (4.93, 0.25) (1.03, 0.18) (1.06, 0.36) (1.63, 0.99) (1.10, 0.40)
Best model generated summary (4.26, 1.01) (4.13, 1.13) (4.67, 0.61) (4.9, 0.30) (2, 1.33) (1.03, 0.18) (2.36, 1.37) (1.16, 0.59)
(b) Human evaluation scores obtained by averaging the ratings provided by three non-clinician on ten conversations
Table A.4: Human evaluation scores. For Completeness, Relevance, Fluency, and Clarity, a rating closer to 5 indicates the best, whereas for Missingness, Hallucination, Contradiction, and Repetition, a rating closer to 1 is preferable.
Annotator 1 Annotator 2
Annotator 1 1.00 0.45
Annotator 2 0.45 1.00
(a) Cohen’s Kappa Coefficient
Annotator 1 Annotator 2
Annotator 1 1.00 0.87
Annotator 2 0.87 1.00
(b) Pearson’s Correlation Coefficient
Table A.5: Inter-rater Reliability (Clinical Annotators)
Annotator 1 Annotator 2 Annotator 3
Annotator 1 1.00 0.70 0.43
Annotator 2 0.70 1.00 0.42
Annotator 3 0.43 0.42 1.00
(a) Cohen’s Kappa Coefficient
Annotator 1 Annotator 2 Annotator 3
Annotator 1 1.00 0.97 0.75
Annotator 2 0.97 1.00 0.75
Annotator 3 0.75 0.75 1.00
(b) Pearson’s Correlation Coefficient
Table A.6: Inter-rater Reliability (Non-Clinical Annotators)
Conversation Generated Summary
Doctor: What is your patient ID? Patient is a 32 year old girl. She is feeling a little
Patient: 1001 tight in her chest and hasn’t been feeling well
Doctor: What is your age? recently. She hasn’t gone to the hospital to see a
Patient: 32 doctor. Her mood is good and she feels mentally
Doctor: What is your gender? tired. She feels tired without doing anything
Patient: Female and doesn’t want to work. She also feels lack
Patient: "Okay" of confidence. She has lost interest in her past
Doctor: "Hello" hobbies.
Doctor: "What are your main problems recently?"
Patient: "I haven’t been feeling well recently, and I feel a little tight in my chest"
Doctor: "Have you ever gone to the hospital to see a doctor?"
Patient: "Not yet, I don’t have much time recently"
Patient: "Maybe it will take two weeks to go"
Doctor: "Hmm, let’s take some time to see if you have any emotional problems recently"
Patient: "There’s nothing wrong with my mood, I just feel mentally tired recently"
Doctor: "Do you feel tired without doing anything?"
Patient: "I feel like this, I don’t want to move"
Doctor: "Then do you feel like you don’t want to work?"
Patient: "I don’t have enough energy to work"
Patient: "Yes"
Patient: "But I have to work"
Patient: "It’s quite stressful"
Doctor: "Have you ever felt that you have lost interest in your past hobbies?"
Patient: "I should still be interested"
Patient: "I just don’t have enough time to develop my hobbies"
Doctor: "Sleep, eat, etc."
Doctor: "Is everything normal?"
Patient: "fairly normal"
Doctor: "Will you feel dizzy or nauseous?"
Patient: "I get dizzy occasionally"
Doctor: "Do you feel lack of confidence? You are always worried about not doing well"
Patient: "No, I don’t have time to worry about this or that"
Doctor: "It sounds like you are doing well lately"
Doctor: "The consultation ends here"
Table A.7: BART-large-CNN generated summary on a sample Chinese psychological conversation taken from Yao et al. (2022) - I
Conversation Generated Summary
Doctor: What is your patient ID? Patient is a 21 year old girl. She is feeling tired
Patient: 1010 and under a lot of pressure to study. Her mood
Doctor: What is your age? is normal and nothing has changed. She does
Patient: 21 not feel unhappy, just tired and tiring easily. Her
Doctor: What is your gender? appetite and sleep is not affected. She feels better
Patient: Female by going for a walk, talking to friends, exercising
Patient: "Hello" and writing.
Patient: "I feel tired recently"
Doctor: "How long has this situation been going on?"
Patient: "About a week or so"
Doctor: "What’s the reason?"
Patient: "Because I am a student, I have been under a lot of pressure to study recently"
Doctor: "Okay, it’s normal to have academic pressure. You can usually find more people to talk to."
Patient: "Okay"
Doctor: "Are you feeling unhappy recently?"
Patient: "No"
Patient: "Just a little tiring easily"
Doctor: "So you can still do the things you were interested in before?"
Patient: "Nothing has changed"
Doctor: "Do you feel inferior easily?"
Patient: "Some people feel that they are always not as good as others"
Doctor: "It’s okay, everyone has their own strengths, don’t envy others"
Patient: "Yeah"
Doctor: "Has there been any change in appetite?"
Patient: "No"
Doctor: "What about sleep?"
Patient: "Neither"
Doctor: "Well, I think you don’t have a big problem. Students are all under academic pressure."
Doctor: "Remember to go out for a walk more often and relax"
Doctor: "Maybe talk to your friends more"
Patient: "Okay, thank you doctor"
Doctor: "Then you can do more things you like"
Doctor: "You can exercise more and write to help relax"
Patient: "Yeah, I will do it"
Doctor: "Then it’s over"
Table A.8: BART-large-CNN generated summary on a sample Chinese psychological conversation taken from Yao et al. (2022) - II

### v2

@@ 3.1 MSE questionnaire design
We identified the absence of a standardized MSE questionnaire and reviewed existing options online. We aimed to create a preliminary version tailored to students, encompassing key components like socialness, mood, attention, memory, frustration tolerance, and social support. This process yielded an 18-question questionnaire. Subsequently, we sought the expertise of clinical psychiatrists to refine the questionnaire further. Their valuable insights were instrumental in vetting the relevance and wording of the questions, resulting in a finalized version of the MSE comprising 12 questions. The final MSE questionnaire is provided in the Appendix LABEL:tab:MSE_Questionnaire.
@@ 3.3 Dialogue representation
We developed a Python script to transform participants’ MSE questionnaire responses into simulated doctor-patient conversations to replicate real-world conversations. This process generated 300 doctor-patient conversation sessions, with 3600 (=12 responses x 300 participants) utterances from participants and an equal number from doctors, totaling 7200 utterances. An anonymized excerpt of such conversation for one participant is presented in Table LABEL:tab:psychiatrist_patient_diaogue in the appendix. Figure 2 shows the average length of utterances for each of the 12 questions. The average length of the dialogue conversation with and without the questionnaire is 3662 and 2054 characters, respectively.
@@ 3.4 Reference human summaries
Furthermore, given the structured nature of the MSE questions, the template was well-suited for summarization purposes. The summary template underwent thorough scrutiny through a rigorous review process involving feedback from three independent reviewers (i.e., graduate researchers). Subsequent revisions were made based on their input, ensuring the summary effectively captured key information while maintaining conciseness, clarity, and correctness. After multiple iterations, the final version of the summary template was approved for use by a psychiatrist, leveraging their domain-specific knowledge. The template utilized for creating summaries for each participant can be found in Appendix LABEL:appendix:_summary_template.
@@ 4 Experiments
The dataset comprising 300 conversations was divided into 200 for training, 50 for validation, and 50 for testing. The Appendix LABEL:appendix:_hyperparameters lists the hyperparameter settings utilized during model training.
@@ 4.1 ROUGE evaluation
We conducted experiments varying the number of epochs for each model to compare their relative performance, as depicted in the Figure LABEL:fig:_ROUGE_metric_with_varying_epochs in the Appendix. This figure showcases the models’ adaptability across different ROUGE metrics as epochs increase. Notably, within just five epochs, the ROUGE-1 score of the BART-large-xsum-samsum model surged from 0.290 to 0.736. Similarly, both the BART-base and BART-large-CNN models demonstrated significant improvements in all ROUGE values within the same timeframe. However, the T5 large model failed to exhibit notable adaptation within five epochs.
@@
To gain insight into the model-generated summaries, we conducted experiments with all models across different numbers of epochs (epochs = 5, 10, 25, 50, 100). After analyzing the output summaries generated by these models, we randomly selected one of the participant’s summaries for further analysis. We found that the pre-trained weights of these models tended to produce incomplete summaries, although they were able to capture smaller contexts of the conversation, as shown in Table LABEL:tab:model_varying_epochs in the Appendix.
@@ 4.2.1 Qualitative findings
Moreover, there was a higher degree of contradiction in the generated summary, although a certain level of contradiction was also observed in the human-generated summary. This discrepancy may arise from either the scenario where the domain expert who created the summary may have missed some interpretation or from expert reviewers having different perceptions. Surprisingly, the model did not exhibit hallucination, which is a major problem in large language models. Furthermore, Table LABEL:tab:_appendix_-_human_evaluator_both_clinical_non_clinical in the Appendix displays the evaluation scores separately from clinicians and non-clinicians. We observed a slight disparity between clinicians and non-clinicians, indicating that clinicians may require a summary with detailed psychological information.
@@ 4.2.2 Inter-rater agreement
We calculated Cohen’s Kappa and Pearson’s correlation coefficient separately for two clinical and three non-clinical annotators (or reviewers). Our clinical annotators achieved a Cohen’s kappa coefficient of 0.45 and a correlation coefficient of 0.87, indicating moderate agreement and strong correlation, respectively. Among non-clinical annotators, annotators 1 and 2 achieved a higher Cohen’s kappa coefficient of 0.7 and a correlation coefficient of 0.97, demonstrating good reliability in their assessments. Table LABEL:tab:Cohen_clinical displays the Cohen’s Kappa coefficient, while Table LABEL:tab:correlation_clinical shows the correlation coefficient among clinical annotators. Similarly, Table LABEL:tab:Cohen_non_clinical presents the Cohen’s Kappa coefficient, and Table LABEL:tab:correlation_non_clinical2 displays the correlation coefficient among non-clinical annotators.
@@ 5 Generalization
Upon reviewing the reviewers’ ratings, we found that the best fine-tuned model’s summary scored well in relevance, fluency, clarity, and repetition as shown in Table 4. However, the generated summary was slightly lacking in terms of missing information, hallucination, and contradiction. Tables LABEL:tab:chinese_dialogue_1 and LABEL:tab:chinese_dialogue_2 in the appendix present dialogue conversations taken from Yao et al. (2022) alongside the corresponding summaries generated by the fine-tuned BART-large-CNN model.